A retailer is considering building a large store. If the local economy experiences expansion, the firm expects the store to earn a $2,000,000 profit next year. If the local economy experiences a contraction, the firm expects the store to lose $400,000 next year. Analysts estimate a 20% chance for the local economy to experience an expansion next year (hence an 80% chance for contraction). What is the expected monetary value (EMV) of building the large store
Answer #1

EMV = 20% * 2000000 + 80% * (-400000) = 400000 - 320000 = $80,000